Marys Boards sells a snowboard, Xpert, that is popular with snowboard enthusiasts. Information relating to Marys purchases
Answer #1

a)

FIFO LIFO
Ending Inventory $2,060 $1,900
Cost of Goods Sold $12,100 $12,260

Notes

Beginning Inventory + Purchases = 141 units

Units sold = 121

Ending Inventory = 141-121 = 20 units

1) In case of FIFO, goods coming in FIRST are sold first. Hence, ending inventory will comprise of inventory purchased at the last. FIFO ending Inventory = 20*103 = $2060

Cost of Goods Sold = Beginning Inventory + Purchases - Ending Inventory

2) FIFO Cost of Goods Sold = 14160-2060 = $12100

3) In case of LIFO, goods coming in LAST are sold first. Hence, ending inventory will comprise of inventory purchased at the beginning. LIFO ending Inventory = 20*95 = $1900

4) LIFO Cost of Goods Sold = 14160-1900 = $12260